OpenStudy (anonymous):

x squared minus 12 times x plus 20 all over x minus 10. x − 2 x minus 2 all over x minus 10 x minus 2 all over 10 x minus 10 all over x minus 2

OpenStudy (zepp):

\[x^2-12x+\frac{20}{x-10}\]

OpenStudy (anonymous):

so when you add those you get an answer?

OpenStudy (zepp):

or \[\large \frac{x^2-12x+20}{x-10}\]?

OpenStudy (anonymous):

yes thats it thanks

OpenStudy (zepp):

Alright so you need to factor the numerator first, do you know how to do it?

OpenStudy (zepp):

At the numator we have \[\large x^2-12x+20\]and want this to be factored Mushroom cheesecake, how do we do that?

OpenStudy (zepp):

It's in fact really simply, you'll need to find two number that's added up give the sum of the leading coefficient (the number sticked to \(x^2\)) and the last constant

OpenStudy (zepp):

and their sum gives the number stucked with x. In our case We need to find two numbers that summed gives 12 and multiplied gives 1 * 20 = 20

OpenStudy (zepp):

gives -12, sorry

OpenStudy (zepp):

These two numbers will be -2 and -10, by find 20's factors. We can rewrite the numerator as \[\large x^2-12x+20=x^2-2x-10x+20\]

OpenStudy (zepp):

And that could be factored into \[x(x-2)-10(x-2)=(x-10)(x-2)\]

OpenStudy (zepp):

Now we got \[\large \frac{(x-10)(x-2)}{(x-10)}=x-2\]

OpenStudy (zepp):

So A) would be the right answer.